# DRS: Deep Question Reformulation With Structured Output

### v1

@@ Abstract
Question answering is a fundamental capability of large language models (LLMs). However, when people encounter completely new knowledge texts, they often ask questions that the text cannot answer due to a lack of understanding of the knowledge. Recent research shows that large language models identify the unanswerability of questions, but they lack the ability to help people reformulate their questions. Even powerful models like GPT-3.5 perform poorly in this regard. To enhance the ability of LLMs to assist humans in reformulating questions to extract relevant knowledge from new documents, we propose a zero-shot method called DRS: Deep Question Reformulation With Structured Output. Our proposed method leverages large language models and the DFS-based algorithm to iteratively search for possible entity combinations and constrain the output with certain entities, effectively improving the capabilities of large language models in this area. Extensive experimental results show that our zero-shot DRS method significantly improves the reformulation accuracy of GPT-3.5 from 23.03% to 70.42% and effectively improves the score of open-source large language models, such as Gemma2-9B, from 26.35% to 56.75%.
@@ 1 Introduction
Question answering is a fundamental capability of large language models (LLMs). In recent years, as the abilities of these models advance significantly from GPT-3 and InstructGPT to GPT-4 Brown et al. (2020); Ouyang et al. (2022); OpenAI et al. (2024), their performance on various question-answering benchmarks improves markedly. Although large language models demonstrate a strong ability to answer a wide range of questions, human knowledge is inherently limited. When people encounter new, unfamiliar texts in specific domains, they often pose questions for which answers cannot be found in the text itself. These are typically referred to as unanswerable questions. Previous research describes this phenomenon as "presupposition errors" and notes that, in real-world scenarios, over 30% of questions cannot be answered due to this issue Gao et al. (2023a); Yu et al. (2023).
@@
Since the advent of BERT and GPT-3 Brown et al. (2020); Devlin et al. (2019), language models have rapidly evolved in question-answering tasks Tan et al. (2023); Wang (2022); Nassiri and Akhloufi (2023); Puri et al. (2020). In recent years, many studies aim to help these models recognize the unanswerability of certain questions, encouraging them to respond appropriately to reduce hallucinations and increase reliability Rajpurkar et al. (2018); Deng et al. (2024); Asai and Choi (2020); Sulem et al. (2021, 2022). However, most of these studies focus solely on making models aware of unanswerable questions, without addressing the modification of unanswerable questions into answerable ones while preserving key entities, which could help users extract valuable information they seek from texts in unfamiliar domains Zhao et al. (2024). More importantly, according to Faustini et al. (2023), in a large-scale industrial experiment, rephrasing unanswerable questions posed to virtual assistants significantly enhances the user experience for millions, which highlights the importance of effectively leveraging LLMs to assist people in question reformulation.
Recent research has begun exploring the ability of large language models to assist in reformulating unanswerable questions. For instance, Zhao et al. (2024) constructed a challenging and high-quality benchmark dataset to evaluate the question-reformulation capabilities of these models. Benchmark tests conducted with well-known models like GPT-3.5 and GPT-4 Ouyang et al. (2022); OpenAI et al. (2024) reveal that, even with the use of Chain of Thought (CoT) reasoning Wei et al. (2023), these models fail to achieve ideal performance in question reformulation. This inspires us to explore effective approaches for leveraging LLMs in question reformulation, enabling people to extract valuable knowledge from unfamiliar texts.
Firstly, to enhance accuracy and efficiency in question reformulation evaluation, we introduce a simple yet effective GPT-4o-mini evaluator as a replacement for the existing Llama2-7B evaluator, which suffers from poor generalization and slow performance. Our GPT-4o-mini evaluator retains the original evaluation criteria while significantly improving accuracy, achieving over 90% compared to the original evaluator’s accuracy of approximately 50%.
Secondly, to enable large language models to effectively assist people in reformulating unanswerable questions from unfamiliar texts, we propose a zero-shot method called DRS: Deep Question Reformulation with Structured Output. This method begins by extracting and filtering entities from the original question, then applies a DFS-based algorithm to identify possible combination of entities, construct text-based statements, and generate corresponding answerable questions - both presented as structured outputs constrained to include selected entities. Finally, the large language model selects the most reliable question from a set of candidates as the final output.
Our method enhances the relevance of reformulated questions to the original by employing a DFS-based algorithm and constraining the model to produce structured outputs, maximizing the likelihood of users obtaining the desired answers. Unlike baseline methods from previous research, our approach takes a more holistic perspective on question generation; the entity constraints ensure greater reliability in the generated questions, resulting in significant accuracy improvements on benchmarks. We conducted extensive evaluations using various LLMs on the dataset proposed by Zhao et al. (2024), which includes texts and questions from diverse sources. The results show that our method significantly boosts various large language models’ reformulation accuracy across all datasets, clearly demonstrating its effectiveness.
In summary, our contributions in this paper are as follows:
@@
(i) We propose a zero-shot method called DRS, which enables LLMs to effectively reformulate unanswerable questions by using a DFS-based combination search to generate structured and reliable outputs.
(ii) Extensive experiments show that zero-shot DRS significantly improves LLMs’ performance, surpassing few-shot CoT by over 100%.
(iii) We introduce the high-speed, high-accuracy GPT-4o-mini evaluator to replace the slow, low-performing Llama2-7B evaluator, enhancing fairness and effectiveness.
@@
Figure 2: The process of our zero-shot DRS method, which mainly contains three different part - entity extraction, dfs combination search with question generation, final candidate selection.
@@ 2 Related Work
Question answering has long been a core focus in natural language processing (NLP) Wang (2022). For many years, researchers have developed new datasets to explore the capabilities of language models in this area, with well-known examples such as CBT, SearchQA, and others Hill et al. (2016); Dunn et al. (2017). These datasets primarily assess a model’s ability to answer questions, without considering the possibility that some questions may be unanswerable.
Since the advent of GPT-3 Brown et al. (2020), the rapid development of large language models has increasingly focused attention on unanswerable questions, whether due to ambiguity or limitations in knowledge Deng et al. (2024); Gao et al. (2023b); Yin et al. (2023). Many studies have proposed methods to address question ambiguities for LLMs, often targeting issues like missing qualifiers that can lead to varied answers Kim et al. (2024); Zhang and Choi (2023).
Extensive research on questions outside the knowledge scope of LLMs has focused on guiding models toward self-reflection, helping them recognize when they lack sufficient knowledge to answer a question and thus reducing hallucinations Yin et al. (2023); Ji et al. (2023); Tonmoy et al. (2024). However, the research often overlooks a common real-life scenario: when people encounter unfamiliar knowledge domains and texts, they may pose questions that appear relevant but are ultimately unanswerable Zhao et al. (2024).
The issue of unanswerable questions in QA systems was first formally addressed in a dataset developed by Rajpurkar et al. (2018). This pioneering work has since been expanded by various studies, such as Yu et al. (2023) analyzing Google search queries and Kim et al. (2023) examining Reddit discussions. Both teams focused on identifying questions with incorrect assumptions or presuppositions that made them impossible to answer accurately based solely on the given document.
Recently, Zhao et al. (2024) advanced research in this field by creating a high-quality dataset of document-related unanswerable questions to explore how LLMs can help people reformulate such questions based on the document. They applied basic approaches, including zero-shot, few-shot, and CoT prompting, to several prominent large language models to evaluate their effectiveness in reformulating unanswerable questions. However, the results were suboptimal. Therefore, in this paper, we propose the zero-shot DRS method to significantly improve large language models’ ability to assist users in question reformulation, further unlocking their potential in this area.
@@ 3 Methodology
Our approach consists of three main steps: (i) Entity extraction and filtering, where large language models are used to extract core entities. (ii) Entity combination search and structured output generation, which employs a DFS-based algorithm to identify entity combinations to form answerable questions. (iii) DFS pruning and candidate selection, where we limit the number of candidate questions and search depth through DFS pruning, then select the most reliable question from the generated candidates for final output. The detailed process of our zero-shot DRS method is illustrated in Figure 2.
@@ 3.1 Entity Extraction and Filtering
For example, when people ask a question like, "When does Rainer Hertrich, the German co-head of EADS, step down?", the important entities are "Rainer Hertrich", "German", and "EADS". These terms represent the key subject of the sentence and the relevant modifiers. Modern large language models can identify important entities from a question. However, these models may also include phrases like "step down", which is a verb phrase, not an important entity, as part of the important entities. This results in the model outputting less precise entities, leading to greater errors in subsequent steps.
@@
(ii) We then apply the large language model again with a zero-shot prompt as an effective classifier to classify the previously extracted entities into five categories: subject, object, predicate, attribute, and others. We retain all entities classified as subject, object, and attribute, as these are the most important components of the question, and discard the others.
[…]
@@
Evaluator QA2 BanditQA BBC Reddit Yelp SQuADv2 Average 200 / 247 300 / 507 50 / 59 100 / 113 50 / 51 300 / 736 1000 / 1713 Llama2-7B Zhao et al. (2024) $39.0039.00$ $68.6768.67$ $14.0014.00$ $79.0079.00$ $44.0044.00$ $72.0072.00$ $52.7852.78$ GPT-4o-mini $90.5090.50$ $92.3392.33$ $88.0088.00$ $92.0092.00$ $90.0090.00$ $91.3391.33$ $90.7090.70$
@@ 3.2 DFS Combination Search and Structured Question Generation
After obtaining reliable entities, we proceed with question reformulation. We use the DFS algorithm to explore combinations of different entities. For any combination that could reformulate a new question, we allow it to generate a completely new question and store it for later selection. Specifically, in this step, we focus on the following processes:
@@
In this step, our algorithm uses a DFS search approach to identify possible entity combinations that can be reformulated into new questions. During the generation process, we find that directly prompting the large language model to reshape the question often results in a low success rate due to the model’s potentially unlimited output content. To address this, we guide the large language model to generate structured outputs by restricting the content required in the generated text. We ensure the reliability of the reformulated question by first generating a statement, then using that statement to create the question. Additionally, to enhance the algorithm’s efficiency, we verify the entities immediately after generating the new question, preventing unnecessary subsequent steps. We also apply pruning to the DFS algorithm by limiting the total number of preselected questions and controlling the search depth.
3.3 Candidate Question Evaluation
In the previous step, we generate a set of new questions through search and recorded the number of entities contained in each. Now, we input these questions along with their respective entity counts into the large language model for reassessment of their answerability. Using a straightforward prompt, we guide the model to evaluate and select the optimal question. When the model determines that multiple questions are answerable based on the document, it selects the question with the highest number of entities and outputs its corresponding label, yielding the final reformulated question.
@@ 4.2 Large Language Models
To effectively evaluate our methodology and compare it with baseline approaches from previous studies, we conduct all experiments using four different LLMs.
@@
Open-Source Models. We also perform experiments using two well-known models developed by Google and Alibaba: Gemma2-9B  [^2] https://huggingface.co/google/gemma-2-9b-it Team et al. (2024) and Qwen2.5-7B  [^3] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct Yang et al. (2024); Team (2024).
@@ 4.4 Implementation Details
In our main experiments using the proposed DRS method, we set the temperature to 0.0 for all large language models to ensure reproducibility. Similarly, for the GPT-4o-mini evaluator, we set the temperature to 0.0 to maintain reliability and consistency in the model’s judgments.
For the two GPT series models, we use the official OpenAI API [^4] https://openai.com/ for inference. For the two open-source models, we deploy them on a single A6000 GPU using bfloat16 precision and use the generate function for text generation. (Additional details are provided in Appendix A).
Model Method    QA2 BanditQA     BBC Reddit     Yelp SQuADv2 Average GPT-3.5 Zero-Shot $28.7428.74$ $11.4111.41$ $13.5613.56$ $18.5818.58$ $19.6119.61$ $36.6936.69$ $21.4321.43$ Zero-Shot w/ CoT $29.1529.15$ $15.6315.63$ $13.5613.56$ $20.3520.35$ $15.6915.69$ $43.7943.79$ $23.0323.03$ Few-Shot $32.7932.79$ $22.6922.69$ $15.2515.25$ $23.0123.01$ $29.4129.41$ $38.4638.46$ $26.9426.94$ Few-Shot w/ CoT $44.9444.94$ $48.7848.78$ $16.9516.95$ $18.5818.58$ $25.4925.49$ $41.2241.22$ $32.6632.66$ Zero-Shot DRS (ours) 81.80 73.20 62.71 75.22 66.67 62.90 70.42 GPT-4o-mini Zero-Shot $36.4436.44$ $15.0815.08$ $15.2515.25$ $23.0123.01$ $15.6915.69$ $47.9347.93$ $25.5725.57$ Zero-Shot w/ CoT $49.3949.39$ $37.5037.50$ $42.3742.37$ $17.7017.70$ $35.2935.29$ $50.1050.10$ $38.7338.73$ Few-Shot $36.4436.44$ $17.8017.80$ $16.9516.95$ $21.2421.24$ $21.5721.57$ $46.5546.55$ $26.7626.76$ Few-Shot w/ CoT $61.1361.13$ $52.8552.85$ $47.4647.46$ $43.3643.36$ $56.8656.86$ $59.5759.57$ $53.5453.54$ Zero-Shot DRS (ours) 88.26 80.16 79.66 83.19 78.43 78.30 81.33 Qwen2.5-7B Zero-Shot $46.1546.15$ $24.0524.05$ $25.4225.42$ $23.8923.89$ $25.4925.49$ $46.7046.70$ $31.9531.95$ Zero-Shot w/ CoT $47.1847.18$ $44.2144.21$ $27.1227.12$ $35.4035.40$ $31.3731.37$ $53.5553.55$ $39.8139.81$ Few-Shot $39.6839.68$ $25.1425.14$ $15.2515.25$ $23.8923.89$ $19.6119.61$ $39.6439.64$ $27.2027.20$ Few-Shot w/ CoT $54.2854.28$ $45.5945.59$ $25.4225.42$ $25.6625.66$ $33.3333.33$ $50.6450.64$ $39.1539.15$ Zero-Shot DRS (ours) 77.73 69.84 55.93 53.10 58.82 60.55 62.66 Gemma2-9B Zero-Shot $36.0336.03$ $21.3321.33$ $20.3420.34$ $18.5818.58$ $19.6119.61$ $42.2142.21$ $26.3526.35$ Zero-Shot w/ CoT $29.1529.15$ $17.1217.12$ $22.0322.03$ $23.0123.01$ $19.6119.61$ $32.9432.94$ $23.9623.96$ Few-Shot $26.3226.32$ $16.1716.17$ $25.4225.42$ $15.0415.04$ $17.6517.65$ $29.3929.39$ $21.6721.67$ Few-Shot w/ CoT $22.6722.67$ $18.0718.07$ $10.1710.17$ $24.7824.78$ $25.4925.49$ $27.4227.42$ $21.4321.43$ Zero-Shot DRS (ours) 59.92 60.73 55.93 59.29 49.02 55.62 56.75

### v3

@@ Abstract
Question answering represents a core capability of large language models (LLMs). However, when individuals encounter unfamiliar knowledge in texts, they often formulate questions that the text itself cannot answer due to insufficient understanding of the underlying information. Recent studies reveal that while LLMs can detect unanswerable questions, they struggle to assist users in reformulating these questions. Even advanced models like GPT-3.5 demonstrate limited effectiveness in this regard. To address this limitation, we propose DRS: Deep Question Reformulation with Structured Output, a novel zero-shot method aimed at enhancing LLMs’ ability to assist users in reformulating questions to extract relevant information from new documents. DRS combines the strengths of LLMs with a DFS-based algorithm to iteratively explore potential entity combinations and constrain outputs using predefined entities. This structured approach significantly enhances the reformulation capabilities of LLMs. Comprehensive experimental evaluations demonstrate that DRS improves the reformulation accuracy of GPT-3.5 from 23.03% to 70.42%, while also enhancing the performance of open-source models, such as Gemma2-9B, from 26.35% to 56.75%.
@@ 1 Introduction
Question answering has emerged as a fundamental capability of large language models (LLMs), with recent advances from GPT-3 and InstructGPT to GPT-4 Brown et al. (2020); Ouyang et al. (2022); OpenAI et al. (2024) demonstrating remarkable improvements on various benchmarks Tan et al. (2023); Wang (2022); Nassiri and Akhloufi (2023). However, when humans encounter unfamiliar knowledge domains, they frequently pose questions that cannot be directly answered from the available text. Recent studies indicate that over 30% of questions in real-world scenarios fall into this category, significantly impacting information access and learning efficiency Gao et al. (2023a); Yu et al. (2023). More importantly, industrial experiments have shown that effectively reformulating such questions can dramatically improve user experience and task completion rates in virtual assistant systems, with potential impact on millions of users Faustini et al. (2023).
@@
A successful question reformulation must satisfy two key criteria: (1) the reformulated question should be answerable based on the given text, and (2) it should preserve the core entities and intent of the original question, ensuring users obtain the information they actually seek. Consider the following example shown in Figure 1: when presented with a text stating "Wasabi is normally consumed in such small quantities that its nutritional value is negligible. The major constituents of raw wasabi root are carbohydrates (23.5%), water (69.1%), fat (0.63%), and protein (4.8%)", users might ask "How many calories are in wasabi?". While this question cannot be directly answered, a well-reformulated version would be "What are the major constituents of raw wasabi root?", which maintains the user’s interest in wasabi’s composition while being answerable from the text.
Previous approaches to handling unanswerable questions broadly fall into three categories: (1) detection methods that focus on identifying unanswerable questions Rajpurkar et al. (2018); Asai and Choi (2020); Sulem et al. (2021, 2022), (2) clarification methods that seek additional information from users Deng et al. (2024); Kim et al. (2023), and (3) reformulation methods that attempt to modify questions into answerable forms Zhao et al. (2024). While recent advances in LLMs have improved performance in question detection and clarification, question reformulation remains challenging due to the difficulty in balancing answerability and intent preservation. Even powerful models like GPT-3.5 achieve only 23.03% accuracy in reformulation tasks, highlighting the significant room for improvement in this area.
To address these challenges, we propose DRS (Deep Question Reformulation with Structured Output), a zero-shot method that enables LLMs to effectively reformulate unanswerable questions. DRS addresses the key challenges through three innovations: (1) a systematic entity-driven approach that ensures intent preservation by explicitly tracking and maintaining key entities from the original question, (2) a structured output framework that ensures answerability by constraining the generation process to incorporate specific entities and generate questions aligned with the corresponding statements derived from the source document, and (3) an efficient DFS-based search strategy, coupled with a candidate question evaluation mechanism, enhances the method’s effectiveness and practicality for real-world applications. Unlike previous methods that often sacrifice one aspect for another, DRS achieves strong performance on both answerability and entity preservation simultaneously.
We conduct extensive experiments on six diverse datasets, comparing DRS with multiple baseline approaches across different types of questions and domains. The results demonstrate that DRS significantly improves reformulation accuracy across all tested LLMs, with GPT-3.5’s performance increasing from 23.03% to 70.42% and open-source model Gemma2-9B improving from 26.35% to 56.75%. Additionally, we introduce a more reliable evaluation framework using GPT-4o-mini, replacing the previous Llama2-7B evaluator to ensure more accurate assessment of reformulation quality.
Our contributions in this paper are threefold:
@@
(i) We propose DRS, a zero-shot method that enables LLMs to effectively reformulate unanswerable questions through entity-driven search and structured outputs.
(ii) We demonstrate through extensive experiments that DRS significantly outperforms existing methods, improving reformulation accuracy by over 100% across various LLMs and datasets.
(iii) We introduce an improved evaluation framework based on GPT-4o-mini, providing more reliable assessment of question reformulation quality.
@@
Figure 2: The complete process of our zero-shot DRS method, which mainly contains three parts - entity extraction and filtering, dfs combination search and structured question generation, and candidate question re-evaluation.
@@ 2 Related Work
Question answering has been a central focus in natural language processing (NLP) Wang (2022); Puri et al. (2020); Nassiri and Akhloufi (2023), with the development of datasets like CBT and SearchQA Hill et al. (2016); Dunn et al. (2017) designed to assess a model’s ability to answer questions. However, these datasets do not address the challenge of unanswerable questions.
With the rise of GPT-3 Brown et al. (2020) and even more powerful large language models (LLMs), the focus has shifted to handling unanswerable questions, which arise due to ambiguity or gaps in knowledge Deng et al. (2024); Gao et al. (2023b); Yin et al. (2023).
Several studies have explored methods for resolving ambiguities in questions, often targeting issues like missing qualifiers that can lead to different interpretations Kim et al. (2024); Zhang and Choi (2023). Many research on questions outside the model’s knowledge scope has focused on guiding models to recognize when they lack sufficient information to answer, thereby reducing hallucinations Yin et al. (2023); Ji et al. (2023); Tonmoy et al. (2024). However, a commonly overlooked real-world scenario involves individuals posing seemingly relevant questions that are unanswerable due to their limited familiarity with the document’s knowledge domain. In such cases, large language models can be employed to reformulate these questions, enabling the inquirers to obtain the information they likely intended to seek.
The issue of document-related unanswerable questions was first systematically addressed in a dataset by Rajpurkar et al. (2018), and later expanded by studies such as Yu et al. (2023), which analyzed Google search queries, and Kim et al. (2023), which examined Reddit discussions. Both studies focused on identifying questions with incorrect assumptions or presuppositions that made them unanswerable.
In recent work, Zhao et al. (2024) introduced a high-quality dataset of document-related unanswerable questions, exploring how large language models can assist in reformulating such questions. They evaluated several well-known large language models using basic approaches like zero-shot and more advanced ones such as few-shot or Chain of Thought (CoT) prompting Wei et al. (2023), but the results were suboptimal. In this paper, we propose the zero-shot DRS method to significantly improve LLMs’ ability to help users reformulate unanswerable questions, advancing progress in this area.
@@ 3 Methodology
Our DRS method combines DFS (Depth-First Search) algorithm with LLMs to address challenges posed by human-generated unanswerable questions involving multiple entities. Directly inputting all entities into LLMs for question reformulation risks overlooking the lack of meaningful relationships between them, leading to unreliable or incoherent answers. To mitigate this, our DFS-based approach systematically explores semantically related entity combinations, ensuring the generation of meaningful questions. Additionally, by controlling the search depth and iteration count, our DRS method strikes a balance between computational efficiency and accuracy, yielding high-quality results with minimal overhead.
Our approach consists of three main steps: (i) Entity extraction and filtering. (ii) DFS combination search and structured question generation. (iii) Candidate question re-evaluation. The detailed process of our zero-shot DRS method is illustrated in Figure 2.
@@ 3.1 Entity Extraction and Filtering
For instance, consider the question: "When does Rainer Hertrich, the German co-head of EADS, step down?". The key entities in this example are "Rainer Hertrich", "German", and "EADS", which represent the core subject and its relevant modifiers. While modern LLMs are capable of extracting important entities from a question, they sometimes include verb phrases such as "step down", which, although significant to the context, do not qualify as entities. This tendency can lead to the inclusion of extraneous elements, resulting in less precise entity identification and subsequently higher error rates in question reformulation task.
To address this limitation, we add a classification framework that categorizes entities into five semantic roles based on their function within the question. This approach facilitates the identification of core entities that align with the user’s intent while filtering out less critical elements, thereby reducing ambiguities in reformulation and enhancing overall precision.
@@
(ii) We then apply the large language model again as an effective entity classifier to classify the previously extracted entities into five categories: subject, object, predicate, attribute, and others. We retain all entities classified as subject, object, and attribute, as these are the most important components of the question, and discard the others.
Subset Name Test Data Size Document Length Question Length Entity Numbers Domain SQuADv2 507 / 1000 $189.74\pm 77.59$ $13.85\pm 4.32$ $2.29\pm 0.95$ Wikipedia QA2 247 / 506 $1109.45\pm 899.83$ $10.29\pm 2.37$ $2.05\pm 0.57$ Mostly Wikipedia BanditQA 736 / 2070 $363.67\pm 202.03$ $10.36\pm 3.16$ $1.70\pm 0.72$ Wikipedia BBC 59 / 278 $652.90\pm 385.85$ $20.61\pm 4.73$ $3.31\pm 1.04$ News Reddit 113 / 313 $569.48\pm 364.64$ $16.58\pm 3.49$ $2.93\pm 0.78$ Social Media Yelp 51 / 165 $486.87\pm 184.44$ $17.72\pm 3.82$ $3.04\pm 0.76$ Review
@@ 3.2 DFS Combination Search and Structured Question Generation
After obtaining reliable entities, we proceed with question reformulation. We use the DFS-based algorithm to explore combinations of different entities. For any combination that could reformulate a new question, we allow it to generate a completely new question and store it for later evaluation. Specifically, in this step, we focus on the following processes:
@@
In this step, our algorithm uses a DFS search approach to identify possible entity combinations that can be reformulated into new questions. During the generation process, we find that directly prompting the large language model to reshape the question often results in a low success rate due to the model’s potentially unlimited output content. To address this, we guide the large language model to generate structured outputs by restricting the entities required in the generated text. We ensure the reliability of the reformulated question by first generating a statement, then using that statement to create the question. Additionally, to enhance the algorithm’s efficiency, we verify the entities immediately after generating the new question, preventing unnecessary subsequent steps. We also apply pruning to the DFS algorithm by limiting the total number of preselected questions and controlling the search depth.
3.3 Candidate Question Re-evaluation
The DFS search process generates multiple candidate questions, each preserving different aspects of the original question. To select the optimal reformulation, we develop a two-stage re-evaluation strategy.
First, we assess each candidate’s answerability by prompting the large language model to verify whether the question can be answered using only the information present in the document. This ensures that our final selection maintains semantic validity and practical utility.
Second, we consider the entity overlap score of each answerable candidate. Questions with higher entity overlap are preferred as they better maintain the user’s original intent. This is quantified by:
Entity Overlap $\text{Entity Overlap Score}=\frac{\lvert\text{Entities}_{\text{cand}}\cap\text{Entities}_{\text{orig}}\rvert}{\lvert\text{Entities}_{\text{orig}}\rvert}$
When multiple candidates achieve similar answerability, we select the one with the highest entity overlap score. As the example illustrated in Figure 2, we identify three questions that are considered potentially answerable by the large language model. Therefore, we allow the model to select the most optimal reformulated question. While both Q2 and Q4 are recognized as clearly answerable, Q4 is ultimately chosen due to its inclusion of a greater number of valid entities. Thus, the model returns Q4 as the final reformulated question to the user. This approach ensures that the final output: (1) can be answered from the given document, (2) maximally preserves the user’s original intent, and (3) maintains semantic coherence with the source text.
@@ 4.2 Large Language Models
To evaluate our methodology and compare it with baseline approaches from previous studies, we conduct all experiments using four different large language models.
@@
Open-Source Models. We also perform experiments using two well-known models developed by Google and Alibaba: Gemma2-9B Team et al. (2024) and Qwen2.5-7B Yang et al. (2024); Team (2024).
@@ 4.3 Metric
Evaluator QA2 BanditQA BBC Reddit Yelp SQuADv2 Average 200 / 247 300 / 507 50 / 59 100 / 113 50 / 51 300 / 736 1000 / 1713 Llama2-7B Zhao et al. (2024) $39.00$ $68.67$ $14.00$ $79.00$ $44.00$ $72.00$ $52.78$ GPT-4o-mini $90.50$ $92.33$ $88.00$ $92.00$ $90.00$ $91.33$ $90.70$
@@ 4.4 Implementation Details
In our main experiments using the proposed DRS method, we set the temperature to 0.0 for all large language models to ensure reproducibility. Similarly, for the GPT-4o-mini evaluator, we set the temperature to 0.0 and top_p to 1.0 to maintain reliability and consistency in the model’s judgments.
For the two GPT series models, we use the official OpenAI API [^2] https://openai.com/ for inference. For the two open-source models, we use the weights released on HuggingFace [^3] https://huggingface.co/ and deploy them on a single NVIDIA RTX A6000 GPU using bfloat16 precision and utilize the generate function for text generation. (Additional details are provided in Appendix A).